What is the probability that the height of a student picked at random has a height greater than 116 cm

1
Expert's answer
2022-02-28T16:13:51-0500

we use z calculations defined as below

z = ( x - u )/ (standard deviation)

where x = 116

u= average height

after which we look for the value of the obtained z score in the normal tables .


we then define

p(x>16) = 1 - P(x<16)

Thus p(x>16) = 1 - ( the value of the obtained z score in the normal table) which becomes the probability that is required.


since in our question we have not been provided with the average height and the standard deviation, then the solution to the problem cannot be obtained.

But given complete data, we use the steps given above to obtain the solution.